# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

### v3

@@ I Introduction
Our paper is organized as follows: In Sec. II, we discuss general bounds for the growth of the norm $||[H,A]^{(k)}||$ for a local operator $A$ and a nearest-neighbor Hamiltonian $H$. We review the result in the general interacting case without localization [30] before discussing the free-fermion case and the case of localization. In Sec. III, we then consider a specific model, the $XXZ$ spin chain. We start with the free-fermion (XX) model without disorder before discussing the Anderson case of quenched disorder. We also consider the Aubry-André model which has a periodic potential incommensurate with the lattice and shows a phase transition at a finite disorder strength which, as we will demonstrate, is clearly visible in the norm growth. We then turn to the interacting Heisenberg (XXX) model and show that results obtained by symbolic computations are inconsistent with localization for all studied disorder strengths. In Sec. IV, we consider Schrieffer-Wolff transformations which can be used to perturbatively transform the Hamiltonian into a basis of quasi-local conserved operators. We show that such a perturbative construction succeeds in the non-interacting case while it will typically fail once interactions are included. The final section contains a summary of our main results and a discussion of the remaining open issues.
II Bounds on operator norm growth
@@
$[H,A]^{(k)}\equiv\underbrace{[H,[H,[...,[H,A]]]]}_{\textrm{k-times}}$ (II.1)
@@
$H=\sum_{I}\mathbbm{1}_{1}\otimes\cdots\otimes\mathbbm{1}_{I-1}\otimes h_{I,I+1}\otimes\mathbbm{1}_{I+2}\otimes\cdots\otimes\mathbbm{1}_{L}$ (II.3)
@@
$A=\sum_{\{\alpha\}}a_{\alpha_{1},\dots,\alpha_{L}}\sigma^{\alpha_{1}}_{1}\otimes\cdots\otimes\sigma^{\alpha_{L}}_{L}\,.$ (II.5)
@@
The two matrix norms which we have defined in Eq. (II) are sub-additive and the $1$-norm is an upper bound for the $2$-norm
@@ II.1 General Case
Let us define $H=\sum_{I}\sum_{a=1}^{C}h_{I}^{a}$ where $h_{I}^{a}$ are local Hamiltonian densities of the form $h_{I}^{a}=J^{a}_{I}\sigma_{I}^{a_{1}}\otimes\sigma_{I+1}^{a_{2}}$. We note that the commutator (II.1) is zero if the local operator $A$ and the local Hamiltonian densities $h^{a}_{I}$ do not form a connected cluster. Let $\{I_{1},\cdots,I_{k}\}$ be a set of points which fulfills the adjacency condition, then
@@
where we have used the sub-additivity property (II). The commutator now is determined by terms which are products of single operators in the $\{\sigma_{i}^{\alpha}\}$ basis and we can use Eq. (II.7) implying that
$||h^{a_{I_{1}}}_{I_{1}}h^{a_{I_{2}}}_{I_{2}}\cdots h^{a_{I_{k}}}_{I_{k}}A||=||h^{a_{I_{1}}}_{I_{1}}||\,||h^{a_{I_{2}}}_{I_{2}}||\cdots||h^{a_{I_{k}}}_{I_{k}}||\,||A||\,.$ (II.9)
@@
$||[H,A]^{(k)}||$$\leq$ $2^{k}\sum_{\{I_{1},\cdots,I_{k}\}}\sum_{\{a_{I}\}}\underbrace{||h||\cdots||h||}_{\text{$k$-many}}\,||A||$ (II.10)
@@
Eq. (II.10) is now a purely combinatorial problem. We need to figure out how many different arrangements of the lattice points $\{I_{1},\cdots,I_{k}\}$ exist such that a connected cluster is formed. Without repeated indices, $j$ local adjacent Hamiltonians form a cluster with $j$ bonds. At each step, one can either attach a local Hamiltonian to the right or to the left of the existing cluster. Thus, as for the random walk in one dimension, there are $2^{j}$ possibilities to form a cluster with $j$ bonds. Lastly, consider the case that we have a cluster of length $j$ and $k\geq j$ local Hamiltonians $h_{I}$. Then the question is how many ways there are to ‘distribute’ those local Hamiltonians on the $j$ bonds of the cluster. This is the same question as asking how many ways there are to partition $k$ elements into $j$ non-empty sets. The answer is given by the Stirling numbers of the second kind, $S(k,j)$. We therefore obtain
$\sum_{\{I_{1},\cdots,I_{k}\}}=\sum_{j=1}^{k}2^{j}S(k,j)=B_{k}(2)$ (II.11)
@@
$B_{k}(2)\sim\frac{\exp[-k(1+\ln W(k/2)-W^{-1}(k/2))]}{\sqrt{W(k/2)+1}}\text{e}^{-2}k^{k}$ (II.13)
@@
This means that the bound (II.12) grows faster than exponentially and almost factorially. We note that the term in the exponential in the second line of (II.1) only becomes negative for $k\gtrsim 20$ thus suppressing the initial almost factorial growth. For large $k$, one can also use that $W(k)\sim\ln k-\ln\ln k$ to see from Eq. (II.13) that $B_{k}(2)\sim(k/\ln k)^{k}$.
@@
To summarize, if a system does not have local operators which are conserved, then the hypothesis is that the bound on the total norm (II.12) is asymptotically tight [30]. Importantly, a norm growth $s(k)\sim(k/\ln k)^{k}$ is not possible in regular graphs or Cayley trees where the norm cannot grow faster than exponential. Many-body systems and the question of localization in such systems are therefore qualitatively different from studies of localization on regular graphs [40, 41]. Next, we consider non-interacting systems which are one example of systems corresponding to regular graphs where stricter, exponential bounds on the norm can be derived.
II.2 Free Fermions
The main difference between a general, interacting system and a free-fermion system is that in the latter case the commutator of a one-body operator with the Hamiltonian always remains a one-body operator. Therefore many of the connected clusters leading to the the bound (II.11) do not contribute. Here, a free-fermion model is a system with a bilinear Hamiltonian of the form $H=\sum_{j,k}h_{jk}c^{\dagger}_{j}c_{k}$ where $c_{j}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $j$. We do not include pair creation and annihilation terms in $H$ but note that the arguments in the following remain valid qualitatively if they are included.
Our starting point to prove a bound in the free-fermion case is the last line of Eq. (II.10). We now want to understand how many distinct connected clusters there are which give a non-zero contribution to the sum. The Hamiltonian in this case is, in general, a sum of $C$ one-body terms, $H=\sum_{I}\sum_{a=1}^{C}h_{I}^{a}$ such as nearest-neighbor hoppings and local potentials. Now consider a connected cluster $\{I\}=\{I_{1},\dots,I_{k}\}$ which gives a non-zero contribution to the sum in Eq. (II.10). Since the commutator $[H,A]^{(k)}$ is always a one-body operator if $A$ is a one-body operator, there is one fermionic operator on a site $i$ and another fermionic operator on a site $j$ for each such cluster. Then, growing the cluster the only contributions which can potentially be non-zero will come from adding a one-body term $h^{a}$ on one of the bonds $I_{i-1},I_{i},I_{j},I_{j+1}$. I.e., there are at most $4$ possibilities to build a new non-zero cluster with $k+1$ elements. While this is, in general, overcounting the number of possibilities we can thus iteratively show that
$\sum_{\{I_{1},\dots,I_{k}\}}\leq 4^{k}$ (II.16)
@@ II.2 Free Fermions
II.3 Localization
In a localized phase, a local operator $A$ is expected to remain localized to a finite region of space of length $\xi_{\rm loc}$ (up to exponentially small tails) when commuted $k$-times with the Hamiltonian. This means that we expect that connected clusters of length $j\gg\xi_{\rm loc}$ give almost no contribution to the norm. Neglecting such clusters leads to
$\sum_{\{I_{1},\cdots,I_{k}\}}\!\!\!\!\!\!\!{\vphantom{\sum}}^{\prime}\;=\sum_{j=1}^{\min(k,\xi_{\rm loc})}2^{j}S(k,j)\,.$ (II.19)
@@ II.3 Localization
$\sum_{j=1}^{\xi_{\rm loc}}2^{j}S(k,j)\sim\sum_{j=1}^{\xi_{\rm loc}}\frac{2^{j}j^{k}}{k!}<\frac{2^{\xi_{\rm loc}}}{(\lceil\xi_{\rm loc}\rceil-1)!}\xi_{\rm loc}^{k}$ (II.20)
@@
where we have used the asymptotic scaling $S(k,j)\sim j^{k}/j!$ of the Stirling numbers of the second kind for $k\gg j$. The operator growth in a localized phase is therefore at most exponential
$||[H,A]^{(k)}||\leq||A||\frac{2^{\xi_{\rm loc}}}{(\lceil\xi_{\rm loc}\rceil-1)!}(2\xi_{\rm loc}CJ)^{k}\,.$ (II.21)
@@
In deriving this bound, we have made the simplifying assumption that the contributions of all clusters with lengths $j\gg\xi_{\rm loc}$ can be neglected. Given that with increasing length also more possibilities exist to distribute the local Hamiltonians $h^{a}_{I}$, one might worry that this assumption is not justified. We will show in App. A that the argument remains valid if clusters with $j>\xi_{\rm loc}$ are included.
We note that both the free-fermion case and the localized case show an exponential growth of the operator norm. Thus, the total operator norm is insufficient to distinguish a non-localized free-fermion phase from a localized interacting or non-interacting phase. The total norm, however, does tell us whether or not a system is ergodic. What does distinguish non-ergodic non-localized from localized phases though is how different terms in the commutator contribute to the norm. If we denote by $[H,A]^{(k)}_{l}$ the terms in the commutator of order $k$ which have support on $l$ lattice sites, then we expect that in the localized case
$s^{(p)}_{l}(k)\equiv||[H,A]^{(k)}_{l}||_{p}\sim\text{e}^{k}\text{e}^{-l}\,.$ (II.22)

### v4

@@ II Outline and main results
Our paper is organized as follows: In Sec. III, we discuss general bounds for the growth of the norm $||[H,A]^{(k)}||$ for a local operator $A$ and a nearest-neighbor Hamiltonian $H$. We review the result in the general interacting case [30] before discussing the free-fermion case. In Sec. IV we then derive bounds for systems where local operators remain strictly exponentially localized and for systems which are unitarily equivalent to an effective Hamiltonian $\tilde{H}$ of local conserved charges. We also discuss the implications of our results for real-time dynamics and transport. In Sec. V, we then consider a specific model, the $XXZ$ spin chain. We start with the free-fermion (XX) model without disorder before discussing the Anderson case of quenched disorder. We also consider the Aubry-André model which has a periodic potential incommensurate with the lattice and shows a phase transition at a finite disorder strength which, as we will demonstrate, is clearly visible in the norm growth. We then turn to the interacting Heisenberg (XXX) model and show that results obtained by symbolic computations are inconsistemt with a strict exponential localization for all studied disorder strengths. In Sec. VI, we consider Schrieffer-Wolff transformations which can be used to perturbatively transform the Hamiltonian into a basis of quasi-local conserved operators. We show that such a perturbative construction succeeds in the non-interacting case while it appears to fail once interactions are included. The final section contains a summary of our main results and a discussion of the remaining open issues.
III General bounds on operator norm growth
@@
$[H,A]^{(k)}\equiv\underbrace{[H,[H,[...,[H,A]]]]}_{\textrm{k-times}}$ (III.1)
@@
$H=\sum_{I}\mathbbm{1}_{1}\otimes\cdots\otimes\mathbbm{1}_{I-1}\otimes h_{I,I+1}\otimes\mathbbm{1}_{I+2}\otimes\cdots\otimes\mathbbm{1}_{L}$ (III.3)
@@
$A=\sum_{\{\alpha\}}a_{\alpha_{1},\dots,\alpha_{L}}\sigma^{\alpha_{1}}_{1}\otimes\cdots\otimes\sigma^{\alpha_{L}}_{L}\,.$ (III.5)
@@
The two matrix norms which we have defined in Eq. (III) are sub-additive and the $1$-norm is an upper bound for the $2$-norm
@@ III.1 General Case
Let us define $H=\sum_{I}\sum_{a=1}^{C}h_{I}^{a}$ where $h_{I}^{a}$ are local Hamiltonian densities of the form $h_{I}^{a}=J^{a}_{I}\sigma_{I}^{a_{1}}\otimes\sigma_{I+1}^{a_{2}}$. We note that the commutator (III.1) is zero if the local operator $A$ and the local Hamiltonian densities $h^{a}_{I}$ do not form a connected cluster. Let $\{I_{1},\cdots,I_{k}\}$ be a set of points which fulfills the adjacency condition, then
@@
where we have used the sub-additivity property (III). The commutator now is determined by terms which are products of single operators in the $\{\sigma_{i}^{\alpha}\}$ basis and we can use Eq. (III.7) implying that
$||h^{a_{I_{1}}}_{I_{1}}h^{a_{I_{2}}}_{I_{2}}\cdots h^{a_{I_{k}}}_{I_{k}}A||=||h^{a_{I_{1}}}_{I_{1}}||\,||h^{a_{I_{2}}}_{I_{2}}||\cdots||h^{a_{I_{k}}}_{I_{k}}||\,||A||\,.$ (III.9)
@@
$||[H,A]^{(k)}||$$\leq$ $2^{k}\sum_{\{I_{1},\cdots,I_{k}\}}\sum_{\{a_{I}\}}\underbrace{||h||\cdots||h||}_{\text{$k$-many}}\,||A||$ (III.10)
@@
Eq. (III.10) is now a purely combinatorial problem. We need to figure out how many different arrangements of the lattice points $\{I_{1},\cdots,I_{k}\}$ exist such that a connected cluster is formed. Without repeated indices, $j$ local adjacent Hamiltonians form a cluster with $j$ bonds. At each step, one can either attach a local Hamiltonian to the right or to the left of the existing cluster. Thus, as for the random walk in one dimension, there are $2^{j}$ possibilities to form a cluster with $j$ bonds. Lastly, consider the case that we have a cluster of length $j$ and $k\geq j$ local Hamiltonians $h_{I}$. Then the question is how many ways there are to ‘distribute’ those local Hamiltonians on the $j$ bonds of the cluster. This is the same question as asking how many ways there are to partition $k$ elements into $j$ non-empty sets. The answer is given by the Stirling numbers of the second kind, $S(k,j)$. We therefore obtain
$\sum_{\{I_{1},\cdots,I_{k}\}}=\sum_{j=1}^{k}2^{j}S(k,j)=B_{k}(2)$ (III.11)
@@
$B_{k}(2)\sim\frac{\exp[-k(1+\ln W(k/2)-W^{-1}(k/2))]}{\sqrt{W(k/2)+1}}\text{e}^{-2}k^{k}$ (III.13)
@@
This means that the bound (III.12) grows faster than exponentially and almost factorially. We note that the term in the exponential in the second line of (III.1) only becomes negative for $k\gtrsim 20$ thus suppressing the initial almost factorial growth. For large $k$, one can also use that $W(k)\sim\ln k-\ln\ln k$ to see from Eq. (III.13) that $B_{k}(2)\sim(k/\ln k)^{k}$.
@@
To summarize, if a system does not have local operators which are conserved, then the hypothesis is that the bound on the total norm (III.12) is asymptotically tight [30]. Importantly, a norm growth $s(k)\sim(k/\ln k)^{k}$ is not possible in regular graphs or Cayley trees where the norm cannot grow faster than exponential. Many-body systems and the question of localization in such systems are therefore qualitatively different from studies of localization on regular graphs [41, 42]. Next, we consider non-interacting systems which are one example of systems which do correspond to regular graphs and where stricter, exponential bounds on the norm can be derived.
III.2 Free Fermions
The main difference between a general, interacting system and a free-fermion system is that in the latter case the commutator of a one-body operator with the Hamiltonian always remains a one-body operator. Therefore many of the connected clusters leading to the the bound (III.11) do not contribute. Here, a free-fermion model is a system with a bilinear Hamiltonian of the form $H=\sum_{j,k}h_{jk}c^{\dagger}_{j}c_{k}$ where $c_{j}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $j$.
Our starting point to prove a bound in the free-fermion case is the last line of Eq. (III.10). We now want to understand how many distinct connected clusters there are which give a non-zero contribution to the sum. The Hamiltonian in this case is, in general, a sum of $C$ one-body terms, $H=\sum_{I}\sum_{a=1}^{C}h_{I}^{a}$ such as nearest-neighbor hoppings and local potentials. Now consider a connected cluster $\{I\}=\{I_{1},\dots,I_{k}\}$ which gives a non-zero contribution to the sum in Eq. (III.10). Since the commutator $[H,A]^{(k)}$ is always a one-body operator if $A$ is a one-body operator, there is one fermionic operator on a site $i$ and another fermionic operator on a site $j$ for each such cluster. Then, growing the cluster the only contributions which can potentially be non-zero will come from adding a one-body term $h^{a}$ on one of the bonds $I_{i-1},I_{i},I_{j},I_{j+1}$. I.e., there are at most $4$ possibilities to build a new non-zero cluster with $k+1$ elements. While this is, in general, overcounting the number of possibilities we can thus iteratively show that
$\sum_{\{I_{1},\dots,I_{k}\}}\leq 4^{k}$ (III.16)
@@ III.2 Free Fermions
IV Localization
Studying the localization of operators when being commuted with the Hamiltonian has the advantage that it is applicable both to the non-interacting and the interacting case. This is in contrast to the localization of wave functions which is only a useful criterion in the non-interacting Anderson case. However, we then have to define precisely what we mean by a local operator staying localized. In the non-interacting Anderson case there is a simple answer: as we will show below, the usual criterion of the single-particle wave function becoming exponentially localized is equivalent to a local operator remaining strictly exponentially localized when commuted with the Hamiltonian,
$\frac{s_{l}(k)}{s_{1}(k)}\sim\exp(-l)\,,$ (IV.1)
with $s_{l}(k)=||[H,A]_{l}^{(k)}||$. As we will also show below, however, strict exponential localization implies that the total operator norm $s(k)\leq\sum_{l}s_{l}(k)$ cannot grow faster than exponential. This does not lead to any contradiction in the Anderson case because we have already shown in Eq. (III.17) that an exponential norm bound always holds in the free-fermion case. It does, however, cause an issue in the many-body case because it is analytically known that the norm always grows faster than exponential in the Ising model with random longitudinal and transverse fields and we will provide strong evidence from symbolic calculations in Sec. V.2.2 that the same is also true in the Heisenberg chain with random fields. I.e., we will demonstrate that a strict exponential localization is not possible in the many-body localized case. If one takes exponential localization of local operators as the defining feature of a localized system, then one has to conclude that many-body localization does not exist.
Another possible definition of many-body localization is, as we have already eluded to earlier, to demand that a unitary transformation exists which maps the microscopic Hamiltonian onto the effective Hamiltonian (I.2). If one could show that the existence of such a transformation implies that local operators in the microscopic model remain exponentially localized when commuted with the Hamiltonian, then one would have proven that many-body localization is also excluded if defined in this way. We are, however, not able to prove a strict exponential bound but rather only the weaker bound (II.4) which does not exclude an almost factorial growth of the total norm. However, even the bound (II.4) predicts an ultimately almost exponential decay of the norm as a function of the length of the operator support if $l\gg k/\kappa$ at fixed order $k$. We will show in Sec. V.2.2 by symbolic computations that there is no evidence for such a weaker type of localization even at very strong disorder.
IV.1 Strict exponential localization
In a localized phase, one might expect that a local operator $A$ remains localized to a finite region of space of length $\xi_{\rm loc}$ (up to exponentially small tails) when commuted $k$-times with the Hamiltonian. We call this strict exponential localization and prove in Sec. V.1.2 that it does hold in the Anderson case.
In any system which is strictly exponentially localized, we expect that connected clusters of length $j\gg\xi_{\rm loc}$ give almost no contribution to the norm. Starting from the general bound (III.10), which is valid both for non-interacting and interacting systems, and neglecting such clusters leads to
$\sum_{\{I_{1},\cdots,I_{k}\}}\!\!\!\!\!\!\!{\vphantom{\sum}}^{\prime}\;=\sum_{j=1}^{\min(k,\xi_{\rm loc})}2^{j}S(k,j)\,.$ (IV.2)
@@ IV.1 Strict exponential localization
$\sum_{j=1}^{\xi_{\rm loc}}2^{j}S(k,j)\sim\sum_{j=1}^{\xi_{\rm loc}}\frac{2^{j}j^{k}}{k!}<\frac{2^{\xi_{\rm loc}}}{(\lceil\xi_{\rm loc}\rceil-1)!}\xi_{\rm loc}^{k}$ (IV.3)
@@
where we have used the asymptotic scaling $S(k,j)\sim j^{k}/j!$ of the Stirling numbers of the second kind for $k\gg j$. The operator growth in a strictly localized phase is therefore at most exponential
$||[H,A]^{(k)}||\leq||A||\frac{2^{\xi_{\rm loc}}}{(\lceil\xi_{\rm loc}\rceil-1)!}(2\xi_{\rm loc}CJ)^{k}\,.$ (IV.4)
@@
In deriving this bound, we have made the simplifying assumption that the contributions of all clusters with lengths $j\gg\xi_{\rm loc}$ can be neglected. Given that with increasing length also more possibilities exist to distribute the local Hamiltonians $h^{a}_{I}$, one might worry that this assumption is too naive. We will show in the Suppl. Mat. [38] that the argument remains valid if clusters with $j>\xi_{\rm loc}$ are included but are assumed to be exponentially suppressed with their length $l$.
Another way to think about strict exponential localization is to take the norm growth of terms in the commutator of order $k$ which have support on $l$ lattice sites
$s^{(p)}_{l}(k)\equiv||[H,A]^{(k)}_{l}||_{p}\sim\text{e}^{k}\text{e}^{-l}$ (IV.5)
directly as the definition of such a phase. Here the exponential decay with $l$ is the strict exponential localization of the operator while the exponential growth with the order $k$ is a consequence of having a finite number of possible clusters with lengths less than $l$ which can be visited before returning to a cluster of length $l$. Along the commutator path the weights are multiplied leading to an $e^{k}$ growth. This is best understood in a graph theoretical language which is discussed in Sec. V. From Eq. (IV.5) it follows immediately that the total norm $s^{(p)}(k)\leq\sum_{l}s^{(p)}_{l}(k)\sim\text{e}^{k}$ because the sum over $l$ converges.